# BGE M3-Embedding: Multi-Lingual, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation

### v3

@@ Abstract
In this paper, we present a new embedding model, called M3-Embedding, which is distinguished for its versatility in Multi-Linguality, Multi-Functionality, and Multi-Granularity. It can support more than 100 working languages, leading to new state-of-the-art performances on multi-lingual and cross-lingual retrieval tasks. It can simultaneously perform the three common retrieval functionalities of embedding model: dense retrieval, multi-vector retrieval, and sparse retrieval, which provides a unified model foundation for real-world IR applications. It is able to process inputs of different granularities, spanning from short sentences to long documents of up to 8192 tokens. The effective training of M3-Embedding involves the following technical contributions. We propose a novel self-knowledge distillation approach, where the relevance scores from different retrieval functionalities can be integrated as the teacher signal to enhance the training quality. We also optimize the batching strategy, enabling a large batch size and high training throughput to ensure the discriminativeness of embeddings. To the best of our knowledge, M3-Embedding is the first embedding model which realizes such a strong versatility. The model and code will be publicly available at https://github.com/FlagOpen/FlagEmbedding.
@@ 1 Introduction
Embedding models are a critical form of DNN application in natural language processing. They encode the textual data in the latent space, where the underlying semantics of the data can be expressed by the output embeddings Reimers and Gurevych (2019); Ni et al. (2021). With the advent of pre-trained language models, the quality of text embeddings have been substantially improved, making them imperative components for information retrieval (IR). One common form of embedding-based IR is dense retrieval, where relevant answers to the query can be retrieved based on the embedding similarity Karpukhin et al. (2020); Xiong et al. (2020); Neelakantan et al. (2022); Wang et al. (2022); Xiao et al. (2023). Besides, the embedding model can also be applied to other IR tasks, such as multi-vector retrieval where the fine-grained relevance between query and document is computed based on the interaction score of multiple embeddings Khattab and Zaharia (2020), and sparse or lexical retrieval where the importance of each term is estimated by its output embedding Gao et al. (2021a); Lin and Ma (2021); Dai and Callan (2020).
@@
The effective training of such a versatile embedding model poses a significant challenge. In our work, the following technical contributions are made to optimize the training quality. Firstly, we propose a novel self-knowledge distillation framework, where the multiple retrieval functionalities can be jointly learned and mutually reinforced. In M3-Embedding, the [CLS] embedding is used for dense retrieval, while embeddings from other tokens are used for sparse retrieval and multi-vector retrieval. Based on the principle of ensemble learning Bühlmann (2012), such heterogenous predictors can be combined as a stronger predictor. Thus, we integrate the relevance scores from different retrieval functions as the teacher signal, which is used to enhance the learning process via knowledge distillation. Secondly, we optimize the batching strategy to achieve a large batch size and high training throughput, which substantially contributes to the discriminativeness of embeddings. Last but not least, we perform comprehensive and high-quality data curation. Our dataset consists of three sources: 1) the extraction of unsupervised data from massive multi-lingual corpora, 2) the integration of closely related supervised data, 3) the synthesization of scarce training data. The three sources of data are complement to each other and applied to different stages of the training process, which lays foundation for the versatile text embeddings.
M3-Embedding exhibits remarkable versatility in our experiments. It achieves superior retrieval quality for a variety of languages, leading to state-of-the-art performances on popular multi-lingual and cross-lingual benchmarks like MIRACL Zhang et al. (2023c) and MKQA Longpre et al. (2021). It effectively learns the three retrieval functionalities, which can not only work individually but also work together for an even stronger retrieval quality. It also well preserves its superior capability across different input granularities within 8192 tokens, which outperforms the existing methods by a notable advantage.
Our work makes the following contributions. 1) We present M3-Embedding, which is the first model which supports multi-linguality, multi-functionality, and multi-granularity. 2) We propose a novel training framework of self-knowledge distillation and efficient batching strategy. And we also perform high-quality curation of training data. 3) Our model, code, and data will all be publicly available, which provides critical resources for both direct usage and future development of text embeddings.
@@ 3 M3-Embedding
M3-Embedding realizes three-fold versatility. It supports a wide variety of languages and handles input data of different granularities. Besides, it unifies the common retrieval functionalities of text embeddings. Formally, given a query in an arbitrary language , it is able to retrieve document in language from the corpus : . In this place, belongs to any of the functions: dense, sparse/lexical, or multi-vector retrieval; can be another language or the same language as .
Data Source |
Language |
Size |
| Unsupervised Data | ||
MTP |
EN, ZH |
291.1M |
S2ORC, Wikipeida |
EN |
48.3M |
xP3, mC4, CC-News |
Multi-Lingual |
488.4M |
NLLB, CCMatrix |
Cross-Lingual |
391.3M |
CodeSearchNet |
Text-Code |
344.1K |
Total |
1.2B |
| Fine-tuning Data | ||
MS MARCO, HotpotQA, NQ, NLI, etc. |
EN |
1.1M |
DuReader, T-Ranking, NLI-zh, etc. |
ZH |
386.6K |
MIRACL, Mr.TyDi |
Multi-Lingual |
88.9K |
MultiLongDoc |
Multi-Lingual |
41.4K |
@@ 3.1 Data Curation
The training of M3-Embedding calls for a large-scale and diverse multi-lingual dataset. In this work, we perform comprehensive data collection from three sources: the unsupervised data from unlabeled corpora, the fine-tuning data from labeled corpora, and the fine-tuning data via synthesization (shown as Table 1). The three data sources complement to each other, which are applied to different stages of the training process. Particularly, the unsupervised data is curated by extracting the rich-semantic structures, e.g., title-body, title-abstract, instruction-output, etc., within a wide variety of multi-lingual corpora, including Wikipedia, S2ORC Lo et al. (2020), xP3 Muennighoff et al. (2022), mC4 Raffel et al. (2019), and CC-News Hamborg et al. (2017). Besides, the well-curated data from MTP Xiao et al. (2023) is directly incorporated. To learn the unified embedding space for cross-lingual semantic matching, the parallel sentences are introduced from two translation datasets, NLLB NLLB Team et al. (2022) and CCMatrix Schwenk et al. (2021). The raw data is filtered to remove potential bad contents and low-relevance samples. In total, it brings in 1.2 billion text pairs of 194 languages and 2655 cross-lingual correspondences.
In addition, we collect relatively small but diverse and high-quality fine-tuning data from labeled corpora. For English, we incorporate eight datasets, including HotpotQA Yang et al. (2018), TriviaQA Joshi et al. (2017), NQ Kwiatkowski et al. (2019), MS MARCO Nguyen et al. (2016), COLIEE Kim et al. (2022), PubMedQA Jin et al. (2019), SQuAD Rajpurkar et al. (2016), and NLI data collected by SimCSE Gao et al. (2021b). For Chinese, we incorporate seven datasets, including DuReader He et al. (2017), mMARCO-ZH Bonifacio et al. (2021), T-Ranking Xie et al. (2023), LawGPT111https://github.com/LiuHC0428/LAW-GPT, CMedQAv2 Zhang et al. (2018), NLI-zh222https://huggingface.co/datasets/shibing624/nli-zh-all, and LeCaRDv2 Li et al. (2023). For other languages, we leverage the training data from Mr. Tydi Zhang et al. (2021b) and MIRACL Zhang et al. (2023c).
Finally, we generate synthetic data to mitigate the shortage of long document retrieval tasks and introduce extra multi-lingual fine-tuning data (denoted as MultiLongDoc). Specifically, we sample lengthy articles from Wiki and MC4 datasets and randomly choose paragraphs from them. Then we use GPT-3.5 to generate questions based on these paragraphs. The generated question and the sampled article constitute a new text pair to the fine-tuning data. Detailed specifications are presented in Appendix A.2.
@@ 3.2 Hybrid Retrieval
M3-Embedding unifies all three common retrieval functionalities of the embedding model, i.e. dense retrieval, lexical (sparse) retrieval, and multi-vector retrieval. The formulations are presented as follows.
@@
Multi-Vector Retrieval. As an extension of dense retrieval, the multi-vector method makes use of the entire output embeddings for the representation of query and passage: , , where is the learnable projection matrix. Following ColBertKhattab and Zaharia (2020), we use late-interaction to compute the fine-grained relevance score: ; and are the lengths of query and passage.
Thanks to the multi-functionality of the embedding model, the retrieval process can be conducted in a hybrid process. First of all, the candidate results can be individually retrieved by each of the methods (the multi-vector method can be exempted from this step due to its heavy cost). Then, the final retrieval result is re-ranked based on the integrated relevance score: .
@@ 3.3 Self-Knowledge Distillation
The training objectives of different retrieval methods can be mutually conflicting with each their. Therefore, the native multi-objective training can be unfavorable to the embedding’s quality. To facilitate the optimization of multiple retrieval functions, we propose to unify the training process on top of self-knowledge distillation. Particularly, based on the principle of ensemble learning Bühlmann (2012), the predictions from different retrieval methods can be integrated as a more accurate relevance score given their heterogeneous nature. In the simplest form, the integration can just be the sum-up of different prediction scores:
@@
The overall training process is a multi-stage workflow (Figure 2). We use an XLM-RoBERTa Conneau et al. (2020) model pre-trained further through the RetroMAE Xiao et al. (2022) method as the base text encoder. Firstly, the text encoder is pre-trained with the massive unsupervised data, where only the dense retrieval is trained in the basic form of contrastive learning. The self-knowledge distillation is applied to the second stage, where the embedding model is fine-tuned to establish the three retrieval functionalities. Both labeled and synthetic data are used in this stage, where hard negative samples are introduced for each query following the ANCE method Xiong et al. (2020). Detailed processing is presented in Appendix B.1.
@@
| Dense | 67.8 | 78.4 | 80.0 | 56.9 | 55.5 | 57.7 | 78.6 | 57.8 | 59.3 | 56.0 | 72.8 | 69.9 | 70.1 | 78.6 | 86.2 | 82.6 | 61.7 | 56.8 | 60.7 |
| Sparse | 53.9 | 67.1 | 68.7 | 43.7 | 38.8 | 45.2 | 65.3 | 35.5 | 48.2 | 48.9 | 56.3 | 61.5 | 44.5 | 57.9 | 79.0 | 70.9 | 36.3 | 32.2 | 70.0 |
| Multi-vec | 69.0 | 79.6 | 81.1 | 59.4 | 57.2 | 58.8 | 80.1 | 59.0 | 61.4 | 58.2 | 74.5 | 71.2 | 71.2 | 79.0 | 87.9 | 83.0 | 62.7 | 57.9 | 60.4 |
| Dense+Sparse | 68.9 | 79.6 | 80.7 | 58.8 | 57.5 | 59.2 | 79.7 | 57.6 | 62.8 | 58.3 | 73.9 | 71.3 | 69.8 | 78.5 | 87.2 | 83.1 | 62.5 | 57.6 | 61.8 |
| All | 70.0 | 80.2 | 81.5 | 59.8 | 59.2 | 60.3 | 80.4 | 60.7 | 63.2 | 59.1 | 75.2 | 72.2 | 71.7 | 79.6 | 88.2 | 83.8 | 63.9 | 59.8 | 61.5 |
@@ 3.4 Efficient Batching
The embedding model needs to learn from diverse and massive multi-lingual data to fully capture the general semantic of different languages. It also needs to keep the batch size as large as possible (where a huge amount of in-batch negatives can be leveraged) so as to ensure the discriminativeness of text embeddings. Given the limitations on GPU’s memory and computation power, people usually truncate the input data into short sequences for high throughput of training and a large batch size. However, the common practice is not a feasible option for M3-Embedding because it needs to learn from both short and long-sequence data to effectively handle the input of different granularities. In our work, we improve the training efficiency by optimizing the batching strategy, which enables high training throughput and large batch sizes.
Particularly, the training data is pre-processed by being grouped by sequence length. When producing a mini-batch, the training instances are sampled from the same group. Due to the similar sequence lengths, it significantly reduces sequence padding (marked in red) and facilitates a more effective utilization of GPUs. Besides, when sampling the training data for different GPUs, the random seed is always fixed, which ensures the load balance and minimizes the waiting time in each training step. Besides, when handling long-sequence training data, the mini-batch is further divided into sub-batches, which takes less memory footprint. We iteratively encode each sub-batch using gradient checkpointing Chen et al. (2016) and gather all generated embeddings. This method can significantly increase the batch size. For example, when processing text with a length of 8192, the batch size can be increased by more than 20 times. For more details please refer to Appendx B.3. Finally, the embeddings from different GPUs are broadcasted, allowing each device to obtain all embeddings in the distributed environment, which notably expands the scale of in-bath negative samples.
However, users may lack sufficient computational resources or data to train a long-text model. Therefore, we also propose an MCLS strategy to enhance the model’s long-text capabilities without the need for training. This strategy leverages multiple CLS tokens to capture text semantics, applied during inference. Refer to Appendix B.2 for more details.

### v4

# M3-Embedding: Multi-Linguality, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation

@@ Abstract
In this paper, we introduce a new embedding model called M3-Embedding, which is distinguished for its versatility in Multi-Linguality, Multi-Functionality, and Multi-Granularity. It provides a uniform support for the semantic retrieval of more than 100 working languages. It can simultaneously accomplish the three common retrieval functionalities: dense retrieval, multi-vector retrieval, and sparse retrieval. Besides, it is also capable of processing inputs of different granularities, spanning from short sentences to long documents of up to 8,192 tokens. The effective training of M3-Embedding presents a series of technical contributions. Notably, we propose a novel self-knowledge distillation approach, where the relevance scores from different retrieval functionalities can be integrated as the teacher signal to enhance the training quality. We also optimize the batching strategy, which enables a large batch size and high training throughput to improve the discriminativeness of embeddings. M3-Embedding exhibits a superior performance in our experiment, leading to new state-of-the-art results on multilingual, cross-lingual, and long-document retrieval benchmarks.111The model, code, and data is publicly available at https://github.com/FlagOpen/FlagEmbedding.
[…]
M3-Embedding: Multi-Linguality, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation
Jianlv Chen♣ Shitao Xiao♠††thanks: Co-first author Peitian Zhang♠ Kun Luo♠ Defu Lian♣∗ Zheng Liu♠††thanks: Corresponding authors ♠ BAAI ♣ University of Science and Technology of China stxiao@baai.ac.cn {namespace.pt,luokun695,zhengliu1026}@gmail.com chenjianlv@mail.ustc.edu.cn liandefu@ustc.edu.cn
@@ 1 Introduction
Embedding models are a critical form of DNN application in natural language processing. They encode the textual data in the latent space, where the underlying semantics of the data can be expressed by the output embeddings Reimers and Gurevych (2019); Ni et al. (2022). With the advent of pre-trained language models, the quality of text embeddings have been substantially improved, making them imperative components for the information retrieval (IR) system. One common form of embedding-based IR application is dense retrieval, where relevant answers to the query can be retrieved based on the embedding similarity Karpukhin et al. (2020); Xiong et al. (2020); Neelakantan et al. (2022); Wang et al. (2022); Xiao et al. (2023). Besides, the embedding model can also be applied to other IR tasks, such as multi-vector retrieval where the fine-grained relevance between query and document is computed based on the interaction score of multiple embeddings Khattab and Zaharia (2020), and sparse or lexical retrieval where the importance of each term is estimated by its output embedding Gao et al. (2021a); Lin and Ma (2021); Dai and Callan (2020).
@@
The training of M3-Embedding poses a significant challenge. In our work, the following technical contributions are made to optimize the embedding quality. Firstly, we propose a novel self knowledge distillation framework, where the multiple retrieval functionalities can be jointly learned and mutually reinforced. In M3-Embedding, the [CLS] embedding is used for dense retrieval, while embeddings from other tokens are used for sparse retrieval and multi-vector retrieval. Based on the principle of ensemble learning Bühlmann (2012), such heterogenous predictors can be combined as a stronger predictor. Thus, we integrate the relevance scores from different retrieval functions as the teacher signal, which is used to enhance the learning process via knowledge distillation. Secondly, we optimize the batching strategy to achieve a large batch size and high training throughput, which substantially contributes to the discriminativeness of embeddings. Last but not least, we perform extensive and high-quality data curation. Our dataset includes three sources: 1) the extraction of unsupervised data from massive multi-lingual corpora, 2) the integration of closely related supervised data, 3) the synthesization of scarce training data. The three data sources are complement to each other and applied to different training stages, which lays a solid foundation for the versatile text embeddings.
M3-Embedding exhibits a remarkable versatility in our experiments. It achieves superior retrieval quality for a variety of languages, leading to state-of-the-art performances on popular multi-lingual and cross-lingual benchmarks like MIRACL Zhang et al. (2023c) and MKQA Longpre et al. (2021). It effectively learns the three retrieval functionalities, which can not only work individually but also work together for an even stronger retrieval quality. It also well maintains its superior capability across different input granularities within 8192 tokens, which outperforms the existing methods by a notable advantage.
Our contributions are summarized as follows. 1) We present M3-Embedding, which achieves unprecedented versatility in multi-linguality, multi-functionality, and multi-granularity. 2) We propose a novel training framework of self-knowledge distillation and optimize the batching strategy for efficient training. We also create high-quality training resource based on comprehensive data curation. 3) Our model, code, and data is publicly available, offering critical resources for both direct usage and future development of text embeddings.
@@ 3 M3-Embedding
M3-Embedding realizes three-fold versatility. It supports a wide variety of languages and handles input data of different granularities. Besides, it unifies the common retrieval functionalities of text embeddings. Formally, given a query in an arbitrary language , it is able to retrieve document in language from the corpus : . In this place, belongs to any of the functions: dense, lexical, or multi-vector retrieval; can be another language or the same language as .
@@ 3.1 Data Curation
M3-Embedding calls for a large-scale and diverse multi-lingual dataset. In this work, we perform comprehensive data collection from three sources: the unsupervised data from unlabeled corpora, the fine-tuning data from labeled corpora, and the fine-tuning data via synthesization (shown as Table 8). The three data sources complement to each other, which are applied to different stages of the training process. Particularly, the unsupervised data is curated by extracting the rich-semantic structures, e.g., title-body, title-abstract, instruction-output, etc., within a wide variety of multi-lingual corpora, including Wikipedia, S2ORC Lo et al. (2020), xP3 Muennighoff et al. (2023), mC4 Raffel et al. (2020), CC-News Hamborg et al. (2017) and the well-curated data from MTP Xiao et al. (2023). To learn the unified embedding space for cross-lingual semantic matching, the parallel sentences are introduced from two translation datasets, NLLB NLLB Team et al. (2022) and CCMatrix Schwenk et al. (2021). The raw data is filtered to remove potential bad contents and low-relevance samples. In total, it brings in 1.2 billion text pairs of 194 languages and 2655 cross-lingual correspondences.
Besides, we collect relatively small but diverse and high-quality fine-tuning data from labeled corpora. For English, we incorporate 8 datasets, including HotpotQA Yang et al. (2018), TriviaQA Joshi et al. (2017), NQ Kwiatkowski et al. (2019), MS MARCO Nguyen et al. (2016), COLIEE Kim et al. (2023), PubMedQA Jin et al. (2019), SQuAD Rajpurkar et al. (2016), and NLI data from SimCSE Gao et al. (2021b). For Chinese, we integrate 7 datasets, including DuReader He et al. (2018), mMARCO-ZH Bonifacio et al. (2021), T2-Ranking Xie et al. (2023), LawGPTLiu et al. (2023), CMedQAv2 Zhang et al. (2018), NLI-zh222https://huggingface.co/datasets/shibing624/nli-zh-all, and LeCaRDv2 Li et al. (2023). For other languages, we leverage the training data from Mr. Tydi Zhang et al. (2021b) and MIRACL Zhang et al. (2023c).
Finally, we generate synthetic data to mitigate the shortage of long document retrieval tasks and introduce extra multi-lingual fine-tuning data (denoted as MultiLongDoc). Specifically, we sample lengthy articles from Wikipedia, Wudao Yuan et al. (2021) and mC4 datasets and randomly choose paragraphs from them. Then we use GPT-3.5 to generate questions based on these paragraphs. The generated question and the sampled article constitute a new text pair to the fine-tuning data. Detailed specifications are presented in Appendix A.2.
@@ 3.2 Hybrid Retrieval
M3-Embedding unifies the common retrieval functionalities of the embedding model, i.e. dense retrieval, lexical (sparse) retrieval, and multi-vector retrieval. The formulation is presented as follows.
@@
Multi-Vector Retrieval. As an extension of dense retrieval, the multi-vector method utilizes the entire output embeddings for the representation of query and passage: , , where is the learnable projection matrix. Following ColBert Khattab and Zaharia (2020), we use late-interaction to compute the fine-grained relevance score: ; and are the lengths of query and passage.
Thanks to the multi-functionality of the embedding model, the retrieval process can be conducted in a hybrid process. First of all, the candidate results can be individually retrieved by each of the methods (the multi-vector method can be exempted from this step due to its heavy cost). Then, the final retrieval result is re-ranked based on the integrated relevance score:
where the values of , and depend on the downstream scenario.
@@ 3.3 Self-Knowledge Distillation
The training objectives of different retrieval methods can be mutually conflicting with each their. Therefore, the native multi-objective training can be unfavorable to the embedding’s quality. To facilitate the optimization of multiple retrieval functions, we propose to unify the training process on top of self-knowledge distillation. Particularly, based on the principle of ensemble learning Bühlmann (2012), the predictions from different retrieval methods can be integrated as a more accurate relevance score given their heterogeneous nature. In the simplest form, the integration can just be the weighted sum of different prediction scores:
Then we compute the weighted sum of , , and as the loss without self-knowledge distillation:
@@
The training process constitutes a multi-stage workflow (Figure 2). In the first place, the text encoder (an XLM-RoBERTa Conneau et al. (2020) model adapted by RetroMAE Xiao et al. (2022) method) is pre-trained with the massive unsupervised data, where only the dense retrieval is trained in the basic form of contrastive learning. The self-knowledge distillation is applied to the second stage, where the embedding model is fine-tuned to establish the three retrieval functionalities. The random initialization of led to poor accuracy and high at the beginning of the training. In order to reduce the impact of this, we set , , , , and during the training process. Both labeled and synthetic data are used in this stage, where hard negative samples are introduced for each query following the ANCE method Xiong et al. (2020). (See Appendix B.1 for more details.)
@@
| Dense | 69.2 | 78.4 | 80.0 | 56.9 | 56.1 | 60.9 | 78.6 | 58.3 | 59.5 | 56.1 | 72.8 | 69.9 | 70.1 | 78.7 | 86.2 | 82.6 | 62.7 | 56.7 | 81.8 |
| Sparse | 53.9 | 67.1 | 68.9 | 43.8 | 38.6 | 45.1 | 65.4 | 35.3 | 48.2 | 48.9 | 56.1 | 61.5 | 44.5 | 57.9 | 79.1 | 70.9 | 36.1 | 32.5 | 70.0 |
| Multi-vec | 70.5 | 79.6 | 81.0 | 59.3 | 57.8 | 62.0 | 80.1 | 59.4 | 61.5 | 58.3 | 74.5 | 71.2 | 71.2 | 79.1 | 87.9 | 83.0 | 63.7 | 58.0 | 82.4 |
| Dense+Sparse | 70.4 | 79.6 | 80.7 | 58.8 | 58.1 | 62.3 | 79.7 | 58.0 | 62.9 | 58.3 | 73.9 | 71.2 | 69.8 | 78.5 | 87.2 | 83.1 | 63.5 | 57.7 | 83.3 |
| All | 71.5 | 80.2 | 81.5 | 59.6 | 59.7 | 63.4 | 80.4 | 61.2 | 63.3 | 59.0 | 75.2 | 72.1 | 71.7 | 79.6 | 88.1 | 83.7 | 64.9 | 59.8 | 83.5 |
@@ 3.4 Efficient Batching
The embedding model needs to learn from diverse and massive multi-lingual data to fully capture the general semantic of different languages. It also needs to keep the batch size as large as possible (introducing a huge amount of in-batch negatives) to ensure the discriminativeness of text embeddings. Given the limitations on GPU’s memory and computation power, people usually truncate the input data into short sequences for high throughput of training and a large batch size. However, the common practice is not a feasible option for M3-Embedding because it needs to learn from both short and long-sequence data to effectively handle the input of different granularities. In our work, we improve the training efficiency by optimizing the batching strategy, which enables high training throughput and large batch sizes.
Particularly, the training data is pre-processed by being grouped by sequence length. When producing a mini-batch, the training instances are sampled from the same group. Due to the similar sequence lengths, it significantly reduces sequence padding (Figure 3, marked in red) and facilitates a more effective utilization of GPUs. Besides, when sampling the training data for different GPUs, the random seed is always fixed, which ensures the load balance and minimizes the waiting time in each training step. Besides, when handling long-sequence training data, the mini-batch is further divided into sub-batches, which takes less memory footprint. We iteratively encode each sub-batch using gradient checkpointing Chen et al. (2016) and gather all generated embeddings. This method can significantly increase the batch size. For example, when processing text with a length of 8192, the batch size can be increased by more than 20 times. (see Appendx B.3 for more details.) Finally, the embeddings from different GPUs are broadcasted, allowing each device to obtain all embeddings in the distributed environment, which notably expands the scale of in-bath negative samples.
For users who are severely limited in computation or data resource, we present an even simpler method called MCLS (Multi-CLS), which simply inserts multiple CLS tokens to the long document during inference, and takes the average of all CLS embeddings as the ultimate embedding of the document. Despite simplicity, it is surprisingly effective in practice. (See Appendix B.2 for more details.)